OpenStudy (anonymous):

Find the circumferenece of the circle given the radius or the diameter: radius=18 centermeters

OpenStudy (anonymous):

How do u do it?

OpenStudy (anonymous):

formula for circumference = 2 * pi * r where r = radius = 18 cms where d = diameter circumference = pi * d

OpenStudy (anonymous):

use the pi button on your calculator or 3.142 will give you a pretty accurate result

OpenStudy (anonymous):

Ok i got it thank you i got 113.04 which is correct in the answer in the back of my math book

OpenStudy (anonymous):

yw